OpenStudy (anonymous):

What is the solution to the system of equations? x + y = 5 x - 2y = 2 I have no idea how to do this lol

OpenStudy (jhannybean):

Use the elimination method to solve for one variable and then use it to find the other.

OpenStudy (midhun.madhu1987):

Subtract the 2 equations and x will be eliminated

OpenStudy (jhannybean):

You've got your two equations, x+y = 5 x-2y = 2 Let's eliminate the x's first and solve for y's. Multiply -1 to the first equation. -1(x+y =5) x - 2y =2 ___________ -x -y = -5 x- 2y = 2 _____________ -3y = -3 y = 1

OpenStudy (midhun.madhu1987):

 using this value of y=1.... find x.